OpenStudy (theedwardsfamily):

Solve for x. x-3/2=-5 A. –4 B. –13 C. –16 D. –7 @iGreen

OpenStudy (igreen):

Is it: \(\dfrac{x-3}{2} = -5\) or is it \(x -\dfrac{3}{2} = -5\)

OpenStudy (theedwardsfamily):

the first one

OpenStudy (eric_d):

Move the 2 to RIGHT HAND SIDE

OpenStudy (eric_d):

Then, add \(3\) on both sides !

OpenStudy (eric_d):

You may give it a try @TheEdwardsFamily

OpenStudy (theedwardsfamily):

ok

OpenStudy (igreen):

Okay, first we multiply 2 to both sides, what do you get?

OpenStudy (theedwardsfamily):

i\how do i multiply 2 on the \(fraction\) side?

OpenStudy (eric_d):

 simplify them